OpenStudy (anonymous):

Calculating tax-equivalent yield. Assume you are in the 35 percent tax bracket and purchase a 4.80 percent tax-exempt municipal bond. Use formula presented in this chapter to calculate the taxable equivalent yield for this investment.

OpenStudy (anonymous):

(1-0.35) x = 4.80% x = 4.80/0.65 = 7.38%